OpenStudy (anonymous):

IM lost and stuck but please tell me how to solve 2.25(4x-4)=-2+10x+12 step by step please =[

OpenStudy (anonymous):

2.25(4x-4)=-2+10x+12 on the left side just distribute, on the right side combine like terms 9x - 9 = 10x + 10 add 9 to both sides 9x = 10x + 19 subtract 9x from both sides 0 = x + 19 subtract 19 from both sides -19 = x

OpenStudy (anonymous):

thank you